OpenStudy (anonymous):

Simplify 125 to the power of negative 2 over 3 ???

OpenStudy (luigi0210):

First what you need to do is flip the number to make it a positive exponent

OpenStudy (anonymous):

so than it is 125 to the power of 3 over2

OpenStudy (anonymous):

??/

OpenStudy (luigi0210):

\[\frac{1}{125^{3/2}}\]

OpenStudy (jdoe0001):

\(\huge a^{\frac{n}{m}} = \sqrt[m]{a^n}\)

OpenStudy (jdoe0001):

or \(\huge a^{-\frac{n}{m}} = \cfrac{1}{\sqrt[m]{a^n}}\)